Project

General

Profile

20160223 » History » Version 13

Blackmore, Victoria, 29 February 2016 11:42

1 1 Boyd, Steven
h1. Ops/Comm Scheduling meeting
2
3 4 Boyd, Steven
Location : CR11 R3 @ 15:30 pm on Tuesday 23rd
4 1 Boyd, Steven
5
Phone access :  ReadyTalk number 
6
US 866-740-1260
7
UK 0800-496-0576
8
Swiss 0800-70-53-51
9
Italy 800182592
10
Access Code: 2155402
11
12
h2. Agenda
13
14 8 Blackmore, Victoria
* *Actions from last meeting*
15
** Pierrick/Mark: Ensure that the pressure and material in the absorber volumes are recorded
16
** Mark: Make more accurate measurements of tracker patch panel helium leak
17
** Pierrick: Progress on monitoring O2 content of tracker with sensors. Monitoring He levels at TOFs?
18
** John G: Resin test progress? (Re: tracker helium leak)
19
** Durga/John N: What is the effect on the beamline of removing GVA1?
20
** Pierrick: Feedback to alarm handler re: corrupted data was showing false positives
21
** Pierrick: Shared services (e.g. water, air) to be included in alarm handler
22
** Durga: Convergence on data quality flags needed
23
** Durga: Progress on testing of reconstruction quality flags
24
25
26 7 Blackmore, Victoria
* DC report - activities in the Hall this week
27 1 Boyd, Steven
* MOM report - Data-taking this week
28
** Lessons learnt during startup
29 5 Blackmore, Victoria
** Data plots from recent shifts
30
** Summary of runs (MOM)
31 2 Boyd, Steven
* Decay Solenoid PS return and calibration
32 9 Blackmore, Victoria
** Kaeser compressor status
33 1 Boyd, Steven
* Focus coil and absorber installation
34 7 Blackmore, Victoria
* Controls and monitoring
35
** Beamstop summary shows beamstop as closed when it's open
36
** Radiation monitor EPICS display freezes for ~2min
37
** Proton absorber switch readout doesn't update
38
** Q5 and Q6 temperature readback still swapped
39
* Scheduling of required hall work and operations
40 10 Blackmore, Victoria
* "List of things to address by February run":http://micewww.pp.rl.ac.uk/projects/operations/wiki/ListOfThingsToAddress
41 1 Boyd, Steven
42
* AOB
43 8 Blackmore, Victoria
** Surveying of TOF1/TOF2/KL/EMR
44 12 Blackmore, Victoria
** Tracker waveguide improvements
45 1 Boyd, Steven
46
* Chair of meetings for next 5 weeks
47 6 Boyd, Steven
** Steve is teaching at this time for the next 4 weeks. DC will chair the meeting.
48 1 Boyd, Steven
49
* DONM : 1/3/2016
50 13 Blackmore, Victoria
51
h2. Minutes
52
53
Present: Paolo, Melissa, Andy N., Paul K., Mike C., Henry, John G., Scott W., Colin, Roy, Alan G., Paul S., Jaroslaw, Yordan, Chris R., Ryan, Ed, Pierrick, Vicky B., Josef, Victoria
54
55
* O2 monitoring on SSD is done (0.5% oxygen in downstream tracker). Monitor sits before the bubbler.  
56
** Plan to take data with one tracker in air and one in helium (Action: Melissa/Ed)
57
** SSU O2 monitoring needs connecting (Action: Craig)
58
* Decision taken to install FC in beamline and evacuate the beampipe on Wed/Thursday
59
* Decay Solenoid
60
** PSU not yet at customs, arrival prob. next week
61
** Mike & Vicky will think about warming up the cold box to check concerns about the compressor
62
* Controls & Monitoring
63
** Action: Pierrick/Paolo, consider increasing the delay before starting EPICS IOC's after a reboot
64
** Q5 and Q6 temperature monitoring are now the correct way around
65
** Action: Perrick, investigate what's wrong with the neutron monitoring IOC (reading pauses for several seconds, then quickly catches up, then pauses, etc.)
66
* Hall scheduling
67
** ISIS need access to turn off a chiller tomorrow morning (John G. organising)
68
** TOF1, TOF2/KL/EMR survey to be arranged (Action: John G)
69
** Beam protection system tests need scheduling (Action: Craig to liase with Mark Arnold)
70
** Camhall03's picture is upside down
71
* AOB
72
** Anyone with an O2 phone should give the number to John G. It can then be added to the signal booster in the hall.
73
** Missing data sections in the archiver (Action: Paolo to talk to Mike C./Vicky B. and find out what's missing and why)